1) Consider point ions of mass 'M' and charge 'e' immersed in a uniform sea of conduction electrons. The ions are imagined to be in stable equilibrium when at regular lattice points. If one ion is displaced a small distance 'r' from its equilibrium position, the restoring force is largely due to the electric charge within the sphere of radius 'r' centered at the equilibrium position. Take the number density of ions (or of conduction electrons)( 3/4)(π)R
1
Expert's answer
2021-09-30T10:19:42-0400

The frequency of an ion can be found by the equation below:


"\omega=\sqrt{\frac{e^2}{4\pi\epsilon_0MR^3}}."